OpenStudy (aaronandyson):

Can someone please explain cross product and dot product rule to me? @Michele_Liano

OpenStudy (aaronandyson):

@Michele_Laino

OpenStudy (michele_laino):

dot product, can also called the \(external\) product, since to each pair of vectors, corresponds a real number, namely an object of different type whereas, the cross product can be called an \(internal\) product, since to each pair of vectors corresponds another vector, namely an object of the same type

OpenStudy (michele_laino):

for example, if we consider the space of geometric vectors, namely each vector is defined by its component: \[{\mathbf{v}} = \left( {x,y,z} \right)\] we can write these formulas

OpenStudy (aaronandyson):

Isnt this related to "determinant"?

OpenStudy (michele_laino):

yes! of course: here is the scalar product: \[{{\mathbf{v}}_1} \cdot {{\mathbf{v}}_2} = {x_1}{x_2} + {y_1}{y_2} + {z_1}{z_2}\]

OpenStudy (michele_laino):

wherein: \[\begin{gathered} {{\mathbf{v}}_1} = \left( {{x_1},{y_1},{z_1}} \right) \hfill \\ {{\mathbf{v}}_2} = \left( {{x_2},{y_2},{z_2}} \right) \hfill \\ \end{gathered} \]

OpenStudy (aaronandyson):

vector of A dot vector of be is abcostheta,right?

OpenStudy (aaronandyson):

b* not be.

OpenStudy (michele_laino):

yes! provided that \(\theta\) is the angle between A and B

OpenStudy (aaronandyson):

then two perpendicular lines will be zero,won't they?

OpenStudy (aaronandyson):

\[î + j = 0\] If there perpendicular?

OpenStudy (aaronandyson):

Because cos 90 is 0.

OpenStudy (michele_laino):

yes! since \(\cos 90=0\)

OpenStudy (michele_laino):

or \(\cos (\pi/2)=0\)

OpenStudy (michele_laino):

now I write the cross product formula, please wait a moment

OpenStudy (aaronandyson):

okay!

OpenStudy (michele_laino):

here is the cross product, please note that the result is another vector: \[\begin{gathered} {{\mathbf{v}}_1} \times {{\mathbf{v}}_2} = \det \left( {\begin{array}{*{20}{c}} {\mathbf{i}}&{\mathbf{j}}&{\mathbf{k}} \\ {{x_1}}&{{y_1}}&{{z_1}} \\ {{x_2}}&{{y_2}}&{{z_2}} \end{array}} \right) = \hfill \\ \hfill \\ = {\mathbf{i}}\left( {{y_1}{z_2} - {y_2}{z_1}} \right) + {\mathbf{j}}\left( {{z_1}{x_2} - {z_2}{x_1}} \right) + {\mathbf{k}}\left( {{x_1}{y_2} - {x_2}{y_1}} \right) \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

wherein, the vectors \[{\mathbf{i}},{\mathbf{j}},{\mathbf{k}}\] are unit vectors along the axes x, y, and z respectively

OpenStudy (aaronandyson):

shouldn't it be -j(z_1x_2 -z_2x_1)

OpenStudy (michele_laino):

 the vector \(v_1 \times v_2\) is a vector orthogonal to the plane defined by the vectors \(v_1\) and \(v_2\)

OpenStudy (michele_laino):

it should be: \[ - {\mathbf{j}}\left( { - {z_1}{x_2} + {z_2}{x_1}} \right)\]

OpenStudy (aaronandyson):

Then are my notes wrong? :/

OpenStudy (aaronandyson):

Ops.sorry I misread the post. There're correct.

OpenStudy (michele_laino):

furthermore, we can write:  \[\left| {{{\mathbf{v}}_1} \times {{\mathbf{v}}_2}} \right| = \left| {{{\mathbf{v}}_1}} \right|\;\left| {{{\mathbf{v}}_2}} \right|\;\sin \alpha ,\quad 0 \leqslant \alpha \leqslant 180\]